OpenStudy (anonymous):

I will give medal!!!! Help please! create three functions in standard form. f(x), with two real rational solutions. g(x), with two real irrational solutions. h(x), with two complex solutions.

OpenStudy (anonymous):

for a polynomial: ax^2 + bx + c = 0 the two solutions/roots would be: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] the parameter which defines what the solutions will be like is:\[\sqrt{b^2 - 4ac}\] if (the discriminant): b^2 - 4ac < 0, the number under the root is negative, which leads to "complex" solutions [containing i = sqrt(-1)] if sqrt(b^2 - 4ac) is rational. so are the two x values. if sqrt(b^2 - 4ac) is irrational. so are the two x values.